# Provable Scaling Laws of Feature Emergence from Learning Dynamics of Grokking

### v4

@@ Appendix C More Experiments
C.1 Use Groups Algorithms Programming (GAP) to get non-Abelian groups
@@
Figure 12: Training modular addition tasks with real weights ($M=23,41,89,127$). Instead of using gradient descent to update the top layer $V$, in every gradient update we use ridge regression solution $V_{\textrm{ridge}}$ with respect to the current $F$ (Eqn. 4). Learning rate is $0.005$, weight decay is 5e55e-5. Number of hidden nodes $K=256$. Test sample is $20\%$ of the full set of $M^{2}$. Using Adam optimizer. Averaged over 55 seeds. The grokking still happens (for $M=23$ check Fig. 13 for completeness). It is slower for $M=23$ but actually faster for $M=41,89,127$, compared to the baseline (Fig. 10).

### v5

@@ Appendix D More Experiments
D.1 Use Groups Algorithms Programming (GAP) to get non-Abelian groups
@@
Figure 12: Training modular addition tasks with real weights ($M=23,41,89,127$). Instead of using gradient descent to update the top layer $V$, in every gradient update we use ridge regression solution $V_{\textrm{ridge}}$ with respect to the current $F$ (Proposition 1). Learning rate is $0.005$, weight decay is 5e55e-5. Number of hidden nodes $K=256$. Test sample is $20\%$ of the full set of $M^{2}$. Using Adam optimizer. Averaged over 55 seeds. The grokking still happens (for $M=23$ check Fig. 13 for completeness). It is slower for $M=23$ but actually faster for $M=41,89,127$, compared to the baseline (Fig. 10).
@@
D.2 Zero-init accelerates the feature learning process
Figure 14: Zero-init accelerates the feature learning process for $M=41$. Check Fig. 16 for the hyperparameters setting. Blue lines are zero-init for $V$ (i.e., $V(0)=0$), red lines are regular initialization.
Figure 15: Zero-init accelerates the feature learning process for $M=89$. Check Fig. 16 for the hyperparameters setting. Blue lines are zero-init for $V$ (i.e., $V(0)=0$), red lines are regular initialization.
Figure 16: Zero-init accelerates the feature learning process for $M=127$. Learning rate is $0.005$. Using Adam optimizer. $\eta=5e-5$. $K=1024$. $\sigma(x)=x^{2}$ and MSE loss. $\text{{Top\_layer\_init\_multiplier}}=0$ means $V(0)=0$, i.e., the top layer is initialized to be 0 (blue lines). If $\text{{Top\_layer\_init\_multiplier}}=1$, then $V$ is initialized regularly (red lines). It is clear that zero-initialization leads to faster grokking (and feature learning) than regular initialization, in particular when data are scarce.
Figure 17: Zero-init accelerates the feature learning process for multi-layer setting. $\text{{other\_layers}}=1$ means adding an additional hidden layer in between top layer $V$ and bottom layer $W$ with residual connection, etc. Learning rate is $0.005$. Using Adam optimizer. $\eta=1e-4$. $K=2048$. $\sigma(x)=\mathrm{ReLU}(x)$ and MSE loss. $\text{{Top\_layer\_init\_multiplier}}=0$ means $V(0)=0$, i.e., the top layer is initialized to be 0 (blue lines). If $\text{{Top\_layer\_init\_multiplier}}=1$, then $V$ is initialized regularly (red lines). It is clear that zero-initialization leads to much faster feature learning than regular initialization in multi-layer settings (often $10\times$).